# MAV: A Multimodal, Multigenre, and Multipurpose Audio-Visual Academic Lecture Dataset

### v1

@@ 5.2.3 Experimental Results
For the objective metrics, we adopt the F0 Frame Error (FFE) Chu and Alwan (2009) and Mel-Cepstral Distortion (MCD) Kubichek (1993). For the subjective metrics, the Mean Opinion Score (MOS) test on a scale of 1 to 5 is performed, which is detailed in Appendix A.4. We randomly choose 20 samples and generate speech segments with the MQTTS model and well-known pre-trained models, including Bark999https://huggingface.co/suno/bark and SpeechT5 Ao et al. (2022)101010https://huggingface.co/microsoft/speecht5_tts. Table 6 shows the experimental results, in which the MQTTS model shows the best performance within all the evaluation metrics. It indicates that the real speech in our dataset can drive AI systems to simulate more natural speech.
@@ 5.3.1 Task Description
The Slide and Script Generation (SSG) Task is designed to promote AI models to understand and reconstruct advanced academic knowledge, thus assisting researchers in creating slides and presentations. The selected paper sentences are provided to supplement rich knowledge. Such models with academic comprehension can facilitate researchers to handle frequently updated academic materials to acquire knowledge and innovation efficiently.
@@ 5.3.2 Benchmark Systems
Having witnessed the excellent performance of large language models (LLMs) and large multimodal models (LMMs) nowadays, we explore their capabilities there. For the "SlideScript" task, we adopt multimodal cascading pipelines and integrated systems. Specifically, the cascading pipelines include fine-tuned LLaMA-2-7B Touvron et al. (2023) and GPT-4 OpenAI (2023a) using slide OCR text and the integrated systems include fine-tuned InstructBLIP-7B Dai et al. (2023) and GPT-4V OpenAI (2023b, c). For the "ScriptSlide" task, we adopt fine-tuned LLaMA-2-7B and GPT-4. For the selection of related paper sentences, We calculate the similarity between the script text and each paper sentence and filter from them. More details can be found in Appendix B.4.
@@ 5.3.3 Experimental Results
| Model | R-1 () | R-2 () | R-L () |
| Slide Script | |||
| LLaMA-2 + OCR | 24.6 | 5.6 | 21.6 |
| GPT-4 + OCR | 26.2 | 5.8 | 23.7 |
| InstructBLIP | 23.7 | 2.2 | 17.7 |
| GPT-4V | 26.4 | 6.2 | 23.7 |
| Script Slide | |||
| LLaMA-2 | 20.3 | 6.2 | 17.3 |
| GPT-4 | 26.8 | 7.7 | 23.0 |
| Range | R-1 () | R-2 () | R-L () |
| Slide Script | |||
| Subset | 24.9 | 6.4 | 21.8 |
| + Paper | 29.2 | 8.7 | 25.8 |
| Script Slide | |||
| Subset | 20.9 | 6.8 | 18.1 |
| + Paper | 22.3 | 7.9 | 19.3 |
ROUGE Lin (2004) scores are reported in Table 7 and Table 8. We get significant observations: 1) The closed-source GPT-4 is stronger than the open-source LLaMA-2 in terms of academic knowledge comprehension. We presume that this is due to a more extensive amount of training data and a larger model size. 2) The latest LMM (GPT-4V) has already exceeded the cascaded pipeline composed of unimodal expert models. It suggests that the LMM not only maintains the ability to process complex textual information but also possesses multi-sensory capabilities, such as the perception and recognition of images and text in slides Yang et al. (2023). We speculate that the reasons for InstructBLIP’s insufficiency are similar to those of LLaMA-2: insufficiently generalised training data and smaller model sizes. 3) Introducing external knowledge can substantially enhance the generation, as shown in Table 8. Quantitative analyses in Appendix C are provided for a more fine-grained view.
As for future development, we are urged to strengthen the open-source model’s ability to understand high-level knowledge and perceive multimodal information. They can be achieved by supplying high-quality data and increasing the model size. We can also mitigate the problem that the open-source multimodal models struggle with processing text in images by increasing the OCR input. Furthermore, how to effectively bring in external knowledge is an essential topic.
@@ 6 Conclusion
We release the Multimodal, Multigenre, and Multipurpose Audio-Visual Dataset with Academic Lectures (MAV) covering a range of academic fields. This dataset contains manually annotated speech transcriptions, slide text, and additional papers, providing a basis for evaluating AI models for recognizing multimodal content and understanding academic knowledge. We detail the creation pipeline and conduct various analyses of the dataset. Furthermore, we build benchmarks and conduct experiments around the dataset. We find there is still large room for existing models to improve perceptions and understanding of academic lecture videos.
@@
Thirdly, we have some deficiencies in mining and labelling for visual information of the dataset. The illustrations in the video frames are not extracted separately like Lee et al. (2023). And, the talking heads of the speakers are not picked up, which can be utilised to research the generation of virtual presenters. In summary, further exploitation of the visual elements in the slides would be a future extension of the dataset.
Finally, the benchmark systems for the (Slide and Script Generation task) SSG task do not involve the generation (for Slide Generation task) and utilization (for Script Generation task) of the visual pictures, instead, we only exploit the text recognised from slides. This is because we can easily extract more accurate information from the text than from visual images. Mining rich information from complex visual frames is regarded as an important research issue in our future work.
@@
This work presents a dataset based on academic speeches with manual annotation in terms of slides and speech. The presentations and talks included in the dataset, are open accessed on the YouTube website. All videos in our dataset are made publicly available by the original creators under the standard YouTube License. Our dataset is provided under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. In terms of the manual annotation, the OCR data is annotated by the team’s researchers, and the speech data annotation, including manual filling, manual combining and TTS scoring is done by the annotators on the Amazon Mechanical Turk. All annotators received clear annotation rules before the annotation and fair payments upon completion. No personal information is requested from any of the annotators.
@@ A.3 Postprocessing for Manual Combination Results
To preserve the consistency of the data after obtaining shown in Figure 4, we perform inverse text normalization111111https://docs.nvidia.com/deeplearning/nemo/user-guide/docs/en/main/nlp/text_normalization/wfst/wfst_text_normalization.html#inverse-text-normalization, add punctuations, and capitalize words using tools121212https://docs.nvidia.com/deeplearning/nemo/user-guide/docs/en/main/nlp/punctuation_and_capitalization.html. Finally, we obtain the word-level timestamps of using Montreal Forced Aligner McAuliffe et al. (2017)131313https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner.
@@ A.4 Manual Annotation Details
The master-granted annotators are assigned to complete manual filling and combination in creating speech transcription. In these two tasks, each worker receives 0.06 USD for each finished sample, respectively. The annotation interfaces shown in Figure 12 and Figure 13 describe the rules as well as the rejection cases of manual filling and combination in detail, respectively. The rejected samples are passed to other workers for annotation.
The experienced researchers in our team are assigned to annotate the OCR data, and the annotation of each sample includes slide deduplication and OCR content correction. Each worker will receive the appropriate remuneration added to their wages. The interface for OCR content correction is shown in Figure 14, where PPOCRLabel141414https://github.com/PaddlePaddle/PaddleOCR/tree/release/2.7/PPOCRLabel is adopted. after completing the annotation, the authors randomly check the annotation results of each person to ensure the quality.
The master-granted annotators are assigned to complete the MOS scoring in the TTS task evaluation. A batch of 6 TTS results from the same audio clip (with the addition of a trap sample used to detect attentiveness). Each worker receives 0.2 USD for each completed batch respectively. Similarly, the annotation interface shown in Figure 15 details the rules as well as the rejection cases, which is inspired by Choi et al. (2023). The rejected samples will be passed to other workers for annotation.
@@ B.1 ASR Benchmark Details
We utilize the CTC-attention model as our ASR baseline, following the instructions of the well-known ESPnet Watanabe et al. (2018) toolkit. More specifically, we follow the espnet2 format due to its more unified and integrated code pattern. We use Conformer Gulati et al. (2020) as our encoder, which is improved from the Transformer by capturing both long-distance information and local representations. Our Conformer encoder consists of 12 blocks, with hidden linear units and attention output size . Attention head is set to 8 and the front CNN kernel size is 31. As for the decoder, we use the Transformer with 6 blocks (,). We use BPE as our recognition unit, which consists of 1k tokens obtained by SentencePiece Kudo and Richardson (2018). In addition to the attention training objectives, We use a certain degree of CTC loss function. CTC weight is 0.3 during training, while 0.2 for infering. SpecAugment Park et al. (2019) is applied with time mask width and frequency mask with . Our model is trained on four 24GB 3090 RTX GPUs for 60 epochs. The top 10 checkpoints after one epoch training are preserved for model averaging. We will release our model and code on ESPnet.
@@ B.2 CASR Benchmark Details
B.2.1 Experimental Setup
Our CASR task is performed under two paradigms. One paradigm is to form the biasing list for an utterance using target biasing words together with distractors that are randomly selected from all rare words, as a simulation of the real-world scenario Aleksic et al. (2015); Zhao et al. (2019); Pundak et al. (2018); Le et al. (2021b). Our experiment uses 1000 distractors to form the biasing list for each utterance, and rare words are defined by removing the top 5k most frequent words from the vocabulary. Another approach is audio-visual CASR Sun et al. (2022) which extracts the biasing list from visual information, such as OCR annotation.
The main structure of our model is an RNN-Transducer Graves (2012), using the Conformer as the encoder, with an intermediate joint network of 320 dimensions. The Conformer has 15 blocks, 1024 hidden linear units, 4 attention heads, and 256 attention output dimensions. After that, a single layer of LSTM is performed. As for TCPGen, we use a 6-layer GCN Kipf and Welling (2017) for future information representation. Following the method proposed in Le et al. (2021a), the biasing lists are extracted by identifying words in the speech transcription of each utterance that also appears in the total rare words list, followed by the addition of plenty of distractors to simulate a real-life setting. The length of the biasing list is set to 500 during training and 1000 during inferencing. Model training lasts for 70 epochs, and the best 10 of the checkpoints are saved for model averaging. Our models are trained on 24GB 3090 RTX GPUs.
Before model training, the rare word selection is first performed. Our speech corpus includes approximately 43k different words in total. We conduct a word frequency analysis on the training set, identifying words with a frequency ranking exceeding 5k as members of the total rare word set. The number of our rare word list is about 38k. We will release the model, code and the rare word list.
@@ B.3 Spontaneous TTS Benchmark Details
We followed the settings recommended by the architecture authors, and their experiments were trained on a subset of GigaSpeech Chen et al. (2021), which is a speech corpus containing transcribed audio from audiobooks, Podcasts, and YouTube. We use the training set of the MAV for MQTTS model training and arbitrarily extract a sufficient number of samples from the Dev and Test Set for speech synthesis to verify the effectiveness of our model. We adopt HiFi-GAN Kong et al. (2020) as the backbone architecture for the quantizer and discriminator, which we trained for 600k steps. After that, a Transformer is trained to autoregressively generate the code representation based on the past code sequence. The last checkpoints from both steps are preserved. It is worth noting that MQTTS training may take a long time. Our models are trained on 24GB 3090 RTX GPUs. We will release our model and code.
@@ B.4 Slide and Script Generation Benchmark Details
For the selection of related paper sentences, we adopt the DistilBERT Sanh et al. (2019) model in the repository of sentence-transformers Reimers and Gurevych (2019) 151515https://huggingface.co/sentence-transformers/msmarco-distilbert-base-v4 to calculate the similarity between the script text and each paper sentence. Then we select the top 3 similarities with similarity greater than 0.5 as the selection results. In the training phase of benchmark systems, we use LoRA Hu et al. (2022) to fine-tune LLaMA-2-7B Touvron et al. (2023). The optimizer we use is AdamW Loshchilov and Hutter (2018). The learning rate curve is 1 round of warm-up and then drops with a peak learning rate of . The total number of training epochs is 3. Our models are trained on 24GB 3090 GPUs and 48GB A6000 GPUs. Fine-tuned InstructBLIP-7B Dai et al. (2023) is another baseline. Specifically, we fine-tune the officially released checkpoint 161616https://storage.googleapis.com/sfr-vision-language-research/LAVIS/models/InstructBLIP/instruct_blip_vicuna7b_trimmed.pth for 20 epochs with the learning rate increasing linearly from to during the first 1000 steps and dropping to 0 as a cosine function. The model is trained on 80GB A100 GPUs with AdamW Loshchilov and Hutter (2018) as the optimizer. The best checkpoints for the SSG task are chosen according to the performance of the development set. As for the generation settings, we use nucleus sampling Holtzman et al. (2020) with a of 0.9. The maximum number of tokens generated is 300. These generation settings are consistent across LLaMA-2, InstructBLIP, GPT-4 and GPT-4V. We will also release our model and code.

### v2

@@ 5.2.3 Experimental Results
For the objective metrics, we adopt the F0 Frame Error (FFE) Chu and Alwan (2009) and Mel-Cepstral Distortion (MCD) Kubichek (1993). For the subjective metrics, the Mean Opinion Score (MOS) test on a scale of 1 to 5 is performed, which is detailed in Appendix A.4. We randomly choose 20 samples and generate speech segments with the MQTTS model and well-known pre-trained models, including Bark171717https://huggingface.co/suno/bark and SpeechT5 Ao et al. (2022)181818https://huggingface.co/microsoft/speecht5_tts. Table 6 shows the experimental results, in which the MQTTS model shows the best performance within all the evaluation metrics. It indicates that the real speech in our dataset can drive AI systems to simulate more natural speech.
@@ 5.3.1 Task Description
| Model | ROUGE-1 () | ROUGE-2 () | ROUGE-L () | BERTScore () | BARTScore () | Human () |
| Slide Script | ||||||
| LLaMA-2+OCR | 24.6 | 5.6 | 21.6 | 0.073 | -4.970 | 3.6 |
| LLaMA-2+OCR | 26.9 | 6.4 | 23.6 | 0.102 | -4.823 | 3.9 |
| GPT-4+OCR | 26.2 | 5.8 | 23.7 | 0.085 | -4.621 | 4.6 |
| InstructBLIP | 24.3 | 2.2 | 17.9 | 0.041 | -5.504 | 2.4 |
| InstructBLIP | 24.7 | 2.3 | 18.2 | 0.046 | -5.514 | 2.5 |
| GPT-4V | 26.4 | 6.2 | 23.7 | 0.092 | -4.630 | 4.7 |
| Script Slide | ||||||
| LLaMA-2 | 20.3 | 6.2 | 17.3 | 0.064 | -6.689 | 3.3 |
| LLaMA-2 | 23.3 | 7.5 | 19.8 | 0.092 | -6.484 | 3.8 |
| GPT-4 | 26.8 | 7.7 | 23.0 | 0.077 | -6.136 | 4.8 |
| Range | ROUGE-1 () | ROUGE-2 () | ROUGE-L () | BERTScore () | BARTScore () | Human () |
| Slide Script | ||||||
| Subset | 24.9 | 6.4 | 21.8 | 0.114 | -5.110 | 3.7 |
| + Paper | 29.2 | 8.7 | 25.8 | 0.161 | -4.806 | 4.1 |
| Script Slide | ||||||
| Subset | 20.9 | 6.8 | 18.1 | 0.061 | -6.737 | 3.3 |
| + Paper | 22.3 | 7.9 | 19.3 | 0.061 | -6.651 | 3.6 |
We define the task formally as follows. Let , and be the slide picture, slide text and the speech transcription of every segment, respectively. For the “SlideScript” task, we need to generate based on . Note that can be viewed as the intermediate result generated from by the OCR model. For the “ScriptSlide” task, we need to generate based on . The Slide and Script Generation (SSG) Task is designed to promote AI models to understand and reconstruct advanced academic knowledge, thus assisting researchers in creating slides and presentations. The selected paper sentences are provided for retrieval-augmented generation (RAG), supplementing rich knowledge. Such models with academic comprehension can facilitate researchers to handle frequently updated academic materials to acquire knowledge and innovation efficiently.
@@ 5.3.2 Benchmark Systems
Having witnessed the excellent performance of large language models (LLMs) and large multimodal models (LMMs) nowadays, we explore their capabilities there. For the "SlideScript" task, we adopt multimodal cascading pipelines and integrated systems. The cascading pipeline consists of an OCR model to extract the text (we directly adopt the OCR annotations) and a language model to generate the answer. Specifically, the cascading pipelines include fine-tuned LLaMA-2 Touvron et al. (2023) and GPT-4 OpenAI (2023a) using slide OCR text and the integrated systems include fine-tuned InstructBLIP Dai et al. (2023) and GPT-4V OpenAI (2023b, c). For the "ScriptSlide" task, we adopt fine-tuned LLaMA-2 and GPT-4. For the selection of related paper sentences, We calculate the similarity between the script text and each paper sentence and filter from them. More details can be found in Appendix B.4.
@@ 5.3.3 Experimental Results
ROUGE Lin (2004), BERTScore Zhang et al. (2019), and BARTScore Yuan et al. (2021) results are reported in Table 7 and Table 8. We also perform manual scoring, where each sample is scored on a scale from 1 (worst) to 5 (best) by 3 annotators on multiple dimensions, including coherence, consistency, fluency and relevance Fabbri et al. (2021). We show their average values, while the detailed values are in Appendix B.5. Scoring is performed on a subset of size 100, detailed in Appendix A.4.
We get significant observations: 1) The open-source models (LLaMA-2, InstructBLIP) show a limited performance improvement when raised from 7B to 13B. Their performances are far from the closed-source models (GPT-4 and GPT-4V). We believe that high-quality pre-training data, e.g., informative corpus and visual QA data which encapsulate multimodal information, is required to enhance their SSG performances beyond just boosting the model size. 2) The latest LMM (GPT-4V) has already exceeded the cascaded pipeline composed of unimodal expert models. It suggests that the LMM not only maintains the ability to process textual information but also possesses multi-sensory capabilities, such as the perception and recognition of the slides Yang et al. (2023). 3) RAG substantially enhances the generation, as shown in Table 8. Qualitative analyses in Appendix C provide a more fine-grained view.
As for future development, we are urged to strengthen the open-source model’s ability to understand high-level knowledge and perceive multimodal information. They can be achieved by increasing the model size and supplying high-quality data. The issue that the open-source multimodal models struggle with recognising text in images can be mitigated by increasing the OCR input. Furthermore, how to effectively bring in external knowledge is an essential topic.
@@ 6 Conclusion
We release the Multimodal, Multigenre, and Multipurpose Audio-Visual Dataset with Academic Lectures (M3AV) covering a range of academic fields. This dataset contains manually annotated speech transcriptions, slide text, and additional extracted papers, providing a basis for evaluating AI models for recognizing multimodal content and understanding academic knowledge. We detail the creation pipeline and conduct various analyses of the dataset. Furthermore, we build benchmarks and conduct experiments around the dataset. We find there is still large room for existing models to improve perceptions and understanding of academic lecture videos.
@@
Thirdly, we have some deficiencies in mining and labelling for visual information of the dataset. The illustrations in the video frames are not extracted separately like Lee et al. (2022, 2023). And, the talking heads of the speakers are not picked up, which can be utilised to research the generation of virtual presenters. In summary, further exploitation of the visual elements in the slides would be a future extension of the dataset.
Finally, the benchmark systems for the Slide Generation task do not involve the generation of visual pictures, instead, we only get the generated slides in the form of text. The generation of slide images requires complex multimodal models, and the corresponding evaluation metrics require more sophisticated designs.
@@
This work presents a dataset based on academic speeches with manual annotation in terms of slides and speech. The presentations and talks included in the dataset, are open accessed on the YouTube website. The copyright remains with the original owners of the videos. Our dataset is provided under a Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License. The annual annotation is done by the team’s researchers and workers on the Amazon Mechanical Turk. All annotators received clear annotation rules before the annotation and fair payments upon completion. No personal information is requested from any of the annotators.
Acknowledgements
This work is supported by National Key R&D Program of China (No. 2022ZD0162101), National Natural Science Foundation of China (No. 62106140) and STCSM (No. 21511101100, No. 22DZ2229005).
@@ A.3 Postprocessing for Manual Combination Results
To preserve the consistency of the data after obtaining shown in Figure 4, we perform inverse text normalization191919https://docs.nvidia.com/deeplearning/nemo/user-guide/docs/en/main/nlp/text_normalization/wfst/wfst_text_normalization.html#inverse-text-normalization, add punctuations, and capitalize words using tools202020https://docs.nvidia.com/deeplearning/nemo/user-guide/docs/en/main/nlp/punctuation_and_capitalization.html. Finally, we obtain the word-level timestamps of using Montreal Forced Aligner McAuliffe et al. (2017)212121https://github.com/MontrealCorpusTools/Montreal-Forced-Aligner.
@@ A.4 Manual Annotation Details
The master-granted annotators are assigned to complete manual filling and combination in creating speech transcription. In these two tasks, each worker receives 0.06 USD for each finished sample, respectively. The annotation interfaces shown in Figure 12 and Figure 13 describe the rules as well as the rejection cases of manual filling and combination in detail, respectively. The failed samples are passed to other workers for annotation.
The experienced researchers in our team are assigned to annotate the OCR data, and the annotation of each sample includes slide deduplication and OCR content correction. Each worker received the appropriate remuneration. The interface for OCR content correction is shown in Figure 14, where PPOCRLabel222222https://github.com/PaddlePaddle/PaddleOCR/tree/release/2.7/PPOCRLabel is adopted. To control the quality of labelling, we sampled the results of 9 annotators individually and checked whether 120 of the annotated images had a high number of missing or incorrect items, with no more than 10 wrong items.
The master-granted annotators are assigned to complete the MOS scoring in the TTS task evaluation. A batch of 6 TTS results from the same audio clip (with the addition of a trap sample used to detect attentiveness). Each worker receives 0.2 USD for each completed batch respectively. Similarly, the annotation interface shown in Figure 15 details the rules as well as the rejection cases, which is inspired by Choi et al. (2023). The failed samples will be passed to other workers for annotation.
As for the manual scoring of SSG tasks, we posted the annotation task on Amazon Mechanical Turk initially but found that the quality was unsatisfactory due to the complexity of the scripts and speech texts. Therefore, three experienced researchers were assigned to carry out the annotation task. Specifically, annotators need to rate the generated slides/scripts on multiple dimensions, including coherence, consistency, fluency and relevance, as referenced in SummEval Fabbri et al. (2021). Each sample is scored on a scale from 1 (worst) to 5 (best) by 3 annotators. Note that we sampled 100 items due to cost and multiple models. The annotation interfaces are shown in Figure 16 and Figure 17.
@@ B.1 ASR Benchmark Details
We construct the Attention-based Encoder-Decoder (AED) with the help of the well-known ESPnet Watanabe et al. (2018) toolkit. We use Conformer Gulati et al. (2020) as the encoder, which consists of 12 blocks, with hidden linear units and attention output size . Attention head is set to 8 and the front CNN kernel size is 31. As for the decoder, we use the Transformer with 6 blocks (,). We use BPE as our recognition unit, which consists of 1k tokens obtained by SentencePiece Kudo and Richardson (2018). In addition to the attention training objectives, We use a certain degree of CTC loss function. CTC weight is 0.3 during training, while 0.2 for infering. SpecAugment Park et al. (2019) is applied with time mask width and frequency mask with . Our model is trained on four 24GB 3090 RTX GPUs for 60 epochs. The top 10 checkpoints after one epoch training are preserved for model averaging.
For the RNN-Transducer (RNN-T), we also adopt Conformer as the encoder, with an intermediate joint network of 320 dimensions. The Conformer has 15 blocks, 1024 hidden linear units, 4 attention heads, and 256 attention output dimensions. After that, a single layer of LSTM is performed.
@@ B.2 CASR Benchmark Details
Our CASR task is performed under two paradigms. One paradigm is to form the biasing list for an utterance using target biasing words together with distractors that are randomly selected from all rare words, as a simulation of the real-world scenario Aleksic et al. (2015); Zhao et al. (2019); Pundak et al. (2018); Le et al. (2021b). Our experiment uses 1000 distractors to form the biasing list for each utterance, and rare words are defined by removing the top 5k most frequent words from the vocabulary. Another paradigm is audio-visual CASR Sun et al. (2022) which extracts the biasing list from visual information, such as OCR annotation.
As for TCPGen, we use a 6-layer GCN Kipf and Welling (2017) for future information representation. Following the method proposed in Le et al. (2021a), the biasing lists are extracted by identifying words in the speech transcription of each utterance that also appears in the total rare words list, followed by the addition of plenty of distractors to simulate a real-life setting. The length of the biasing list is set to 500 during training and 1000 during inferencing. Model training lasts for 70 epochs, and the best 10 of the checkpoints are saved for model averaging. Our models are trained on 24GB 3090 RTX GPUs.
Before model training, the rare word selection is first performed. Our speech corpus includes approximately 43k different words in total. We conduct a word frequency analysis on the training set, identifying words with a frequency ranking exceeding 5k as members of the total rare word set. The number of our rare word list is about 38k.
@@ B.3 Spontaneous TTS Benchmark Details
We followed the settings recommended by the architecture authors, and their experiments were trained on a subset of GigaSpeech Chen et al. (2021), which is a speech corpus containing transcribed audio from audiobooks, Podcasts, and YouTube. We use the training set of the M3AV for MQTTS model training and arbitrarily extract a sufficient number of samples from the Dev and Test Set for speech synthesis to verify the effectiveness of our model. We adopt HiFi-GAN Kong et al. (2020) as the backbone architecture for the quantizer and discriminator, which we trained for 600k steps. After that, a Transformer is trained to autoregressively generate the code representation based on the past code sequence. The last checkpoints from both steps are preserved. It is worth noting that MQTTS training may take a long time. Our models are trained on 24GB 3090 RTX GPUs.
@@ B.4 Slide and Script Generation Benchmark Details
For the selection of related paper sentences, we adopt the DistilBERT Sanh et al. (2019) model in the repository of sentence-transformers Reimers and Gurevych (2019) 232323https://huggingface.co/sentence-transformers/msmarco-distilbert-base-v4 to calculate the similarity between the script text and each paper sentence. Then we select the top 3 candidates with similarity greater than 0.5 as the selection results. In the training phase of benchmark systems, we use LoRA Hu et al. (2022) to fine-tune LLaMA-2 Touvron et al. (2023). The optimizer we use is AdamW Loshchilov and Hutter (2018). The learning rate curve is 1 epoch of warm-up and then drops with a peak learning rate of . The total number of training epochs is 3. Our models are trained on 24GB 3090 GPUs and 48GB A6000 GPUs. Fine-tuned InstructBLIP Dai et al. (2023) is another baseline. Specifically, we fine-tune the officially released checkpoint for 20 epochs with the learning rate increasing linearly from to during the first 1000 steps and dropping to 0 as a cosine function. The model is trained on 80GB A100 GPUs with AdamW Loshchilov and Hutter (2018) as the optimizer. The best checkpoints are chosen according to the performance of the development set. As for the generation settings, we use nucleus sampling Holtzman et al. (2020) with a of 0.9. The maximum number of tokens generated is 300. These generation settings are consistent across LLaMA-2, InstructBLIP, GPT-4 and GPT-4V.
B.5 Detailed Manual Scoring Results on SSG Tasks
We show the detailed scoring results in Table 9 and Table 10, which include the coherence (Coh.), consistency (Cons.), fluency (Flu.), relevance (Rel.) and average value (Avg.) between the hypothesis and its reference text.
| Model | Coh. () | Cons. () | Flu. () | Rel. () | Avg. () |
| Slide Script | |||||
| LLaMA-2+OCR | 3.2 | 3.0 | 4.2 | 4.1 | 3.6 |
| LLaMA-2+OCR | 3.7 | 3.4 | 4.1 | 4.3 | 3.9 |
| GPT-4+OCR | 4.0 | 4.4 | 5.0 | 5.0 | 4.6 |
| InstructBLIP | 1.6 | 2.0 | 2.6 | 3.2 | 2.4 |
| InstructBLIP | 2.0 | 2.1 | 2.6 | 3.3 | 2.5 |
| GPT-4V | 4.2 | 4.5 | 5.0 | 5.0 | 4.7 |
| Script Slide | |||||
| LLaMA-2 | 3.1 | 3.4 | 3.4 | 3.3 | 3.3 |
| LLaMA-2 | 3.5 | 3.6 | 3.9 | 4.0 | 3.8 |
| GPT-4 | 4.6 | 4.4 | 5.0 | 5.0 | 4.8 |
| Range | Coh. () | Cons. () | Flu. () | Rel. () | Avg. () |
| Slide Script | |||||
| Subset | 3.4 | 3.2 | 4.2 | 4.1 | 3.7 |
| + Paper | 4.1 | 4.2 | 4.1 | 4.1 | 4.1 |
| Script Slide | |||||
| Subset | 3.2 | 3.4 | 3.3 | 3.3 | 3.3 |
| + Paper | 3.6 | 3.7 | 3.5 | 3.4 | 3.6 |